# Topological phases of extended Su-Schrieffer-Heeger-Hubbard model

### v1

@@ IV SUMMARY AND DISCUSSION
When the system is at quarter-filling, long-range AFM order emerges, leading to the disappearance of long-range spin correlation and a significant increase of the entanglement entropy. Interestingly, if an external magnetic field is applied at this case, the system reverts to a quasi-half-filling state under the magnetic field, reintroducing long-range spin correlation and reducing entanglement entropy to near zero.
@@ Appendix C Spin correlation and entropy base on Exact Diagonalization
Figure 9: ED results same as Fig 4. (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=0.7$, topological edge states appear at [$1_{st},3_{rd},6_{th},8_{th}$] sites. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0$ and there is a long-range anti-ferromagnetic order. (c) Same as (a) with $\Delta_{J_{3}-J_{1}}=-0.7$, and topological edge states appear at [$1_{st},8_{th}$] sites. And $\omega$ is the degeneracy of ground state. (d) The entanglement Rényi entropy between subsystem [$3_{rd},6_{th}$] sites and the remaining sites. (e) Same as (d) but subsystem [$1_{st},8_{th}$] sites and (f) with subsystem [$1_{st},3_{rd},6_{th},8_{th}$] sites. In (d) $\sim$ (f), the position where the transition occurs is far from $\Delta_{J_{3}-J_{1}}=0$, which is result of finite-size effects.
@@
Figure 11: Same as Fig 7 for $J_{1}=J_{3}=0$, and $U=10$ for (a) $\sim$ (e). (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{33}-J_{11}}=-0.75$. (b) Same as (a) with $\Delta_{J_{33}-J_{11}}=0.75$ and topological edge states appear at [$2_{nd},(N-1)_{th}$] sites. (c) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{33}-J_{11}}=-0.75$ above the critical magnetic field. (d) Same as (c) but $\Delta_{J_{33}-J_{11}}=0.75$. (e) Rényi entanglement entropy with changing external magnetic field and hopping amplitude difference $\Delta_{J_{33}-J_{11}}$. (f) Critical field strength at various of the on-site interaction and hopping amplitude difference and it can be seen that for a certain $\Delta_{J_{33}-J_{11}}$, the value of the critical magnetic field will decrease as U increases.

### v2

@@ IV SUMMARY AND DISCUSSION
When the system is at quarter-filling, long-range AFM order emerges, leading to the disappearance of long-range spin correlation and a significant increase of the entanglement entropy. Interestingly, if an external magnetic field is applied at this case, the system reverts to a quasi-half-filling state, reintroducing long-range spin correlation and reducing entanglement entropy to near zero.
@@ Appendix A Comparison of ED and CP-AFQMC results with 1D $N=8$ chain and a puzzle
Figure 9: ED results same as Fig 4. (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{3}-J_{1}}=0.7$, topological edge states appear at [$1_{st},3_{rd},6_{th},8_{th}$] sites. (b) Same as (a) with $\Delta_{J_{3}-J_{1}}=0$ and there is a long-range anti-ferromagnetic order. (c) Same as (a) with $\Delta_{J_{3}-J_{1}}=-0.7$, and topological edge states appear at [$1_{st},8_{th}$] sites. And $\omega$ is the degeneracy of ground state. (d) The entanglement Rényi entropy between subsystem [$3_{rd},6_{th}$] sites and the remaining sites. (e) Same as (d) but subsystem [$1_{st},8_{th}$] sites and (f) with subsystem [$1_{st},3_{rd},6_{th},8_{th}$] sites. In (d)-(f), the position where the transition occurs is far from $\Delta_{J_{3}-J_{1}}=0$, which is result of finite-size effects.
Figure 11: Same as Fig 7 for $J_{1}=J_{3}=0$, and $U=10$ for (a)-(e). (a) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{33}-J_{11}}=-0.75$. (b) Same as (a) with $\Delta_{J_{33}-J_{11}}=0.75$ and topological edge states appear at [$2_{nd},(N-1)_{th}$] sites. (c) Spin correlation $\langle S_{z,j}S_{z,k}\rangle$ with $\Delta_{J_{33}-J_{11}}=-0.75$ above the critical magnetic field. (d) Same as (c) but $\Delta_{J_{33}-J_{11}}=0.75$. (e) Rényi entanglement entropy with changing external magnetic field and hopping amplitude difference $\Delta_{J_{33}-J_{11}}$. (f) Critical field strength at various of the on-site interaction and hopping amplitude difference and it can be seen that for a certain $\Delta_{J_{33}-J_{11}}$, the value of the critical magnetic field will decrease as U increases.